The total revenue fun

tion (TR), in terms of P and the

ompany's quantity, if a

ommodity sells for R199.00

are?

[1℄ T R = 6 000 − 30P; Q = 30.

[2℄ T R = 6 000P − 30P

2

; Q = 30.

[3℄ T R = P(6 000 − 30P); Q = 24.

[4℄ T R = 6 000 − 30P; Q = 24.


1
Expert's answer
2021-07-19T10:55:39-0400

"Q=6000-30P=6000-30\cdot199=30"

"TR=QP=P(6000-30P)"


Answer: [2] T R = 6 000P − 30P2; Q = 30.